OpenStudy (anonymous):

11.) Refer to right triangle ACB. If m Ang A = 30° and BC = 5, what is the length of AC? Ang C=90 5 sqrt(2) 5 5 sqrt(3) 10

OpenStudy (anonymous):

Alright given the angles: A=30 B=60 *(180-90-30)* C=90 Given the sides: a=5 Use the Law of Sines: Lengh AC also known as b b= 5/sin(30 (sin(60) b= 8.66

OpenStudy (anonymous):

Sorry, but still don't understand. So, the lengh of AC =8.66. Should that # be rounded because that # is not in the answers?

OpenStudy (anonymous):

Let me double check something...Did the problem give you the degrees for Angle A and C? If so then according to the Law of Sines length of AC would be 8.66 or something withing reasonable range...like 9.

OpenStudy (anonymous):

Angle A =30 and angle C= 90

OpenStudy (anonymous):

Yea answer is ~8.66